# Probing quantum entanglement with Generalized Parton Distributions at the Electron-Ion Collider

### v2

@@ II.2 Transverse photon
$\gamma^{+}=\frac{1}{\sqrt{2}}\left(\gamma^{0}+\sin\theta\gamma^{r}+\cos\theta\gamma^{k}\right),\qquad\gamma^{1}=-\cos\theta\gamma^{r}+\sin\theta\gamma^{k},\qquad\gamma^{2}=\gamma^{n}.$ (32)
@@
where in the first inequality we maximized in $k_{\perp}$. The last inequality is saturated when $X=Y$ and X=YX=-Y, which occurs at $z=0$ and $\bar{z}=0$, respectively. Since the colliding particles $e+p$ or $\gamma+p$ are unpolarized, (34) can be interpreted as single spin asymmetry (SSA), analogous to the transverse polarization of hyperons in unpolarized hadron collisions Bunce:1976yb; Lundberg:1989hw; HERMES:2007fpi; COMPASS:2021bws. [^3] The transverse polarization of final state particles can also arise if initial state particles are transversely polarized. See Cheng:2025zaw for a recent study at the EIC. As is well known, SSA requires a phase, and therefore it cannot be generated in tree-level calculations in parity-conserving theories like QCD. Theoretical descriptions usually invoke higher-twist distribution and/or fragmentation functions, or their transverse momentum dependent generalizations Liang:1997rt; Kanazawa:2000cx; Anselmino:2001js; Zhou:2008fb; Koike:2017fxr; Kang:2021kpt. However, a phase can also be generated in perturbation theory beyond the tree level Dharmaratna:1996xd; Benic:2024fvk, and this is what happens here. Since GPD amplitudes involve a loop, they are naturally complex. (Notice the $i\epsilon$ terms in (10).) (34) clearly shows that the nonvanishing polarization results from the interference between their real and imaginary parts. In contrast, in the Regge limit, $X,Y$ and $I_{L}^{g}$ become purely imaginary, see the next subsection. This is why no polarization was found in Fucilla:2025kit.
@@
Unlike in the longitudinal case, the $C^{T}$ matrix depends on the nonperturbative structure of the proton through their dependence on the quark and gluon GPDs.
One might have expected that $B^{T}_{n}(z)=\bar{B}^{T}_{n}(z)$ and $C^{T}_{rk}(z)=C^{T}_{kr}(z)$ from CP symmetry, as in the case of $q\bar{q}\to t\bar{t}$ and $gg\to t\bar{t}$ productions in QCD Bernreuther:2015yna. However, neither of these relations hold in the present problem since the initial state $\gamma+{\mathbb{P}}$ is asymmetric (see also Qi:2025onf). Instead, the following relations hold exactly
@@
For the valence quarks which we treat as massless $m_{u,d}=0$, this is trivial. For massive quarks $q=s,c,b$, (41) is a consequence of the property $F_{q}(x,\xi)=-F_{q}(-x,\xi)$ of the nonvalence quark GPDs which results in the relations
@@
II.3 Regge limit
Let us quickly check consistency with the result of Fucilla:2025kit obtained in the Regge limit $\xi\to 0$. In this limit, the quark exchange contributions $I_{L}^{q},I_{T}^{q_{1}},I_{T}^{q_{2}}$ are subdominant and neglected altogether. The gluon exchange contributions $I_{L}^{g},I_{T}^{g}$ are dominantly imaginary, coming from the $i\epsilon$ prescription in the integrals (10). Keeping only the most singular term in the imaginary part, one finds Braun:2005rg
@@ II.3 Regge limit
where it is assumed that the gluon GPD has a Regge behavior $H_{g}(\xi,\xi)\sim 1/\xi^{\alpha}$ with $\alpha\ll 1$. On the other hand, the result of Fucilla:2025kit is expressed in terms of the following two integrals
@@
where $T$ is the T-matrix of the $q\bar{q}$ pair (‘color dipole’) scattering off the target proton. To make a connection with the collinear GPD calculation, we expand the integrand in powers of the intrinsic transverse momentum $\bm{p}$} and use the relation between $T(p)$ and $H_{g}(\xi,\xi)$ Hatta:2017cte. We find
@@
Inserting these relations into $A^{T/L}$, and noting that $X\approx(\bar{z}-z)I_{T}^{g}$, $Y\approx I_{T}^{g}$, one recovers the dijet cross section originally calculated in the $k_{\perp}$-factorization framework Nikolaev:1994cd; Bartels:1996ne, as already noted in Braun:2005rg. Furthermore, inserting into (36)-(II.2), we find full agreement with Fucilla:2025kit. Note that $T_{1}$ linearly vanishes as $k_{\perp}\to\mu$ Fucilla:2025kit. Consistently, $I_{L}^{g}$ has the same behavior in the present approximation.
III Entanglement, Bell nonlocality and magic
@@
III.1 Entanglement
According to the Peres-Hordecki criterion Peres:1996dw; Horodecki:1997vt, for generic two-qubit systems, the necessary and sufficient condition for separability (i.e., no entanglement) is that the partial transpose of the density matrix
@@ III.1 Entanglement
is nonnegative definite. (The symbol $\cal T$} denotes ‘transpose.’) In the present problem with $\rho=\rho^{L/T}$, it is convenient to use the spin eigenstates along the direction $\hat{n}$. In this basis, the density matrix is block diagonal Afik:2020onf
@@
In Fucilla:2025kit, it has been noticed that in the Regge limit,
@@
In the longitudinal case, the $q\bar{q}$ pair is always maximally entangled because $C_{nn}^{L}=1$ and $\Delta_{1}^{L}=2>0$. In the transverse case where $\Delta_{2}^{T}=-2C_{nn}^{T}\geq 0$, the pair is almost always entangled except when $C_{nn}^{T}=0$, which occurs in certain kinematical limits such as $k_{\perp}=0$ or $k_{\perp}=\infty$. (52) is a direct consequence of the following relations obeyed by the components of $C^{L/T}$
@@
Away from the Regge limit, (53) continue to hold in the longitudinal case. However, in the transverse case we find from (36)-(II.2)
@@
(56) is an example of the so-called ‘Bell diagonal’ states Lang:2010xtw which are mixed states comprised only of the Bell states
@@
with $\sum_{i}p_{i}=1$. It is known that a Bell diagonal state is separable if $p_{i}\leq\frac{1}{2}$ for all $i$. In the present problem, the pair is separable only if $|X|=|Y|$ where $C_{nn}^{T}=0$. If either $|X|$ or $|Y|$ (accidentally) vanishes, then one of the Bell states is realized. Note that in the Regge limit, $X\approx(1-2z)I_{T}^{g}$ and $Y\approx I_{T}^{g}$. Thus the state is separable when $z=0,1$, and becomes the Bell state $|\Phi^{-}\rangle$ at $z=\frac{1}{2}$.
III.2 Bell nonlocality
In terms of the correlation matrix $C$, the Bell-CHSH inequality reads Bell:1964kc; Clauser:1969ny
@@ III.2 Bell nonlocality
where $\vec{n}_{i}$ ($i=1,2,3,4$) are unit vectors $|\vec{n}_{i}|=1$. Depending on the matrix $C_{ab}$, quantum two-qubit states can violate this inequality for deliberate choices of $\{\vec{n}_{i}\}$. When this happens, we say the $q\bar{q}$ pair exhibits ‘Bell nonlocality.’ The necessary and sufficient condition for Bell nonlocality is that the largest two of the three eigenvalues $\mu_{3}\leq\mu_{2}\leq\mu_{1}$ of the matrix $C^{\cal T}C$ satisfy Horodecki:1995nsk
@@
Thus the inequality is always maximally violated in the longitudinal case $C_{nn}^{L}=1$, while in the transverse case it is violated if $C_{nn}^{T}\neq 0$. Since $C_{nn}^{T}\leq 0$, this happens to be the same condition as (50), (52) for nonvanishing entanglement Fucilla:2025kit. In other words, Bell nonlocality and entanglement are equivalent. This is remarkable because, in general, the states that exhibit Bell nonlocality form a subset of the states that are entangled. Away from the Regge limit, the Bell-CHSH inequality continues to be maximally violated in the longitudinal case. In the transverse case, it is always violated for the $q=u,d$ quarks because (53) also holds for massless quarks. For massive quarks, there will be kinematical regions where the inequality is not violated. On general grounds, we expect that this region is broader than the region where the pair is separable (not entangled).
@@
III.3 Magic
Finally, let us consider another measure of entanglement called ‘magic’. (See White:2024nuc for an introduction friendly to high energy physicists.) It is known that entanglement alone does not guarantee quantum advantage Gottesman:1998hu. Entangled pairs with nonvanishing magic are crucial for quantum computers to outperform classical ones, and there has been interest in the high energy physics community whether collider experiments can produce such pairs White:2024nuc; Liu:2025qfl; Gargalionis:2025iqs. To quantify magic, we adopt the ‘stabilizer Rényi entropy’ Leone:2021rzd; White:2024nuc which takes the following form for the present problem
@@ III.3 Magic
using the same basis as in (48). (The sign combination is arbitrary, except that the 14 and 41 components are complex conjugate to each other.) (66) is significantly smaller than the theoretical upper bound for pure two-qubit states Liu:2025frx
@@
However, we point out that this limit can be saturated only by states with nonzero polarization vectors, namely, $B,\bar{B}\neq 0$ in the generic parametrization (22). Since $B^{L}=\bar{B}^{L}=0$, the maximal value (69) cannot be expected. A better comparison may be the range of $M_{2}$ for the subset of states with vanishing polarization. In Appendix A, we revisit the derivation of Liu:2025frx by imposing an extra constraint $B=\bar{B}=0$ and obtain a stronger bound
@@
which is still larger than (66) by a factor of about 2. In Appendix A, we further show that (66) is the maximum value of $M_{2}$ if one restricts to an even smaller subset of states. We note that the numbers $\ln\frac{9}{5}$ and $\ln\frac{4}{3}$ have been previously encountered in Liu:2025qfl as the maximal amount of magic that can be generated from certain initial stabilizer states in QED processes.
IV Numerical results
In this section we delineate kinematical regions in the parameter space $(z,k_{\perp})$ where the $q\bar{q}$ pair exhibits entanglement and Bell nonlocality. We also compute the stabilizer Rényi entropy $M_{2}^{T}$ for the transversely polarized photon. We employ the Goloskokov-Kroll model Goloskokov:2006hr; Goloskokov:2007nt of the quark and gluon GPDs. This model has been fitted to reproduce the exclusive vector meson production data at HERA. At small skewness, the gluon GPD features the Regge behavior
@@ IV Numerical results
with $Q_{0}^{2}=4\,{\rm GeV}^{2}$ and $\mu_{h}=Q$ in DIS Goloskokov:2006hr; Goloskokov:2007nt). Thus the model should produce results consistent with Fucilla:2025kit in the Regge limit. Adapting to our process, we use $\mu_{h}={\rm max}\{Q,m,k_{\perp}\}$. The quark masses are fixed as $m_{u}=m_{d}=0$, $m_{s}=0.093$ GeV, $m_{c}=1.27$ GeV and $m_{b}=4.18$ GeV. For light quark production $q=u,d,s$, we include both the quark and gluon GPD contributions. For heavy quark production $q=c,b$, we include only the gluon GPD contribution (i.e., $I^{q}_{L/T}=0$ in this case) since the $c,b$-quark GPDs are unavailable (and expected to be negligible). Using the GK model, we numerically evaluate the integrals (10) on grid points on a $50\times 50$ lattice in the $(z,k_{\perp})$ plane
@@
and smoothly interpolate the results. $k_{\perp}^{\rm max}$ is chosen appropriately for a given value of the $\gamma+p$ CM energy $W$. The integrals can be evaluated in a numerically stable way. For very small values of $\xi$, this is achieved by using the contour deformation technique as explained around (43) of Hatta:2025vhs.
In DIS at $Q\neq 0$, the measured cross section (6) is the linear combination of the transverse and longitudinal contributions. Unfortunately, it is difficult to experimentally separate the two contributions at the EIC since typically $\varepsilon\approx 1$ at high energy colliders (see however Klest:2025yik). This means that the density matrices must be averaged

### v3

@@ III.2 Transverse photon
$\gamma^{1}=-\cos\theta\gamma^{r}+\sin\theta\gamma^{k},\qquad\gamma^{2}=\gamma^{n}.$ (37)
@@
where in the first inequality we maximized in $k_{\perp}$. The last inequality is saturated when $X=Y$ and X=YX=-Y, which occurs at $z=0$ and $\bar{z}=0$, respectively. Since the colliding particles $e+p$ or $\gamma+p$ are unpolarized, (39) can be interpreted as single spin asymmetry (SSA), analogous to the transverse polarization of hyperons in unpolarized hadron collisions Bunce et al. (1976); Lundberg et al. (1989); Airapetian et al. (2007); Alexeev et al. (2022). [^4] The transverse polarization of final state particles can also arise if initial state particles are transversely polarized. See Cheng et al. (2025) for a recent study at the EIC. As is well known, SSA requires a phase, and therefore it cannot be generated in tree-level calculations in parity-conserving theories like QCD. Theoretical descriptions usually invoke higher-twist distribution and/or fragmentation functions, or their transverse momentum dependent generalizations Liang and Boros (1997); Kanazawa and Koike (2001); Anselmino et al. (2002); Zhou et al. (2008); Koike et al. (2017); Kang et al. (2022). However, a phase can also be generated in perturbation theory beyond the tree level Dharmaratna and Goldstein (1996); Benić et al. (2024), and this is what happens here. Since the scattering amplitudes (10) are obtained by convoluting GPDs with a hard factor, they are naturally complex. (Notice the $i\epsilon$ terms in (10).) The equation (39) clearly shows that the nonvanishing polarization results from the interference between their real and imaginary parts. In contrast, in the Regge limit, $X,Y$ and $I_{L}^{g}$ become purely imaginary, see the next subsection. This is why no polarization was found in Fucilla and Hatta (2026).
@@
Unlike in the longitudinal case (29), the $C^{T}$ matrix depends on the nonperturbative structure of the proton through their dependence on the quark and gluon GPDs. Importantly, the resulting transverse density matrix $\rho^{T}$ in general represents a mixed state ${\rm Tr}[(\rho^{T})^{2}]<1$ Fucilla and Hatta (2026). This is due to the combined effect of the transverse photon polarization sum and finite longitudinal momentum transfer $2\xi P^{-}$ from the proton to the $q\bar{q}$ pair in the form of GPD convolutions (${\bf\Delta}\approx 0$ in the present approximation). On the other hand, the proton spin does not affect the entanglement property of the pair because we sum over the initial and final proton spins and there is no spin flip during scattering. The density matrix would take a similar form if the target were a spinless pion, with the pion GPDs replacing the proton GPDs.
One might have expected that $B^{T}_{n}(z)=\bar{B}^{T}_{n}(z)$ and $C^{T}_{rk}(z)=C^{T}_{kr}(z)$ from CP symmetry, as in the case of $q\bar{q}\to t\bar{t}$ and $gg\to t\bar{t}$ productions in QCD Bernreuther et al. (2015). However, neither of these relations hold in the present problem since the initial state $\gamma+{\mathbb{P}}$ is asymmetric (see also Qi et al. (2025)). Instead, the following relations hold exactly
@@
(Note that $1-2z\to-(1-2z)$ under the transformation $z\to\bar{z}=1-z$.) For the valence quarks which we treat as massless $m_{u,d}=0$, this is trivial. For massive quarks $q=s,c,b$, (46) is a consequence of the property $F_{q}(x,\xi)=-F_{q}(-x,\xi)$ of the nonvalence quark GPDs which results in the relations
@@
III.3 Regge limit
Let us quickly check consistency with the result of Fucilla and Hatta (2026) obtained in the Regge limit $\xi\to 0$. In this limit, the quark exchange contributions $I_{L}^{q},I_{T}^{q_{1}},I_{T}^{q_{2}}$ are subdominant and neglected altogether. The gluon exchange contributions $I_{L}^{g},I_{T}^{g}$ are dominantly imaginary, coming from the $i\epsilon$ prescription in the integrals (10). Keeping only the most singular term in the imaginary part, one finds Braun and Ivanov (2005)
@@ III.3 Regge limit
where it is assumed that the gluon GPD has a Regge behavior $H_{g}(\xi,\xi)\sim 1/\xi^{\alpha}$ with $\alpha\ll 1$. On the other hand, the result of Fucilla and Hatta (2026) is expressed in terms of the following two integrals
@@
where $T$ is the T-matrix of the $q\bar{q}$ pair (‘color dipole’) scattering off the target proton. To make a connection with the collinear GPD calculation, we expand the integrand in powers of the intrinsic transverse momentum $\bm{p}$} and use the relation between $T(p)$ and $H_{g}(\xi,\xi)$ Hatta et al. (2017). We find
@@
Inserting these relations into $A^{T/L}$, and noting that $X\approx(\bar{z}-z)I_{T}^{g}$, $Y\approx I_{T}^{g}$, one recovers the dijet cross section originally calculated in the $k_{\perp}$-factorization framework Nikolaev and Zakharov (1994); Bartels et al. (1996), as already noted in Braun and Ivanov (2005). Furthermore, inserting into (41)-(III.2), we find full agreement with Fucilla and Hatta (2026). Note that $T_{1}$ linearly vanishes as $k_{\perp}\to\mu$ Fucilla and Hatta (2026). Consistently, $I_{L}^{g}$ has the same behavior in the present approximation.
IV Entanglement, Bell nonlocality and magic
@@
IV.1 Entanglement
According to the Peres-Hordecki criterion Peres (1996); Horodecki (1997), for generic two-qubit systems, the necessary and sufficient condition for separability (i.e., no entanglement) is that the partial transpose of the density matrix
@@ IV.1 Entanglement
is nonnegative definite. (The symbol $\cal T$} denotes ‘transpose.’) In the present problem with $\rho=\rho^{L/T}$, it is convenient to formally introduce the spin eigenstates $|\pm\rangle_{n}=\frac{1}{\sqrt{2}}(|+\rangle_{k}\pm|-\rangle_{k})$ along the direction $\hat{n}$. In the basis $\{|++\rangle_{n},|+-\rangle_{n},|-+\rangle_{n},|--\rangle_{n}\}$, the (transposed) density matrix is block diagonal Afik and de Nova (2021)
@@
Note that nonnegativity of the density matrix ensured by the condition (55) is not affected by unitary Wigner rotations. This is the physical reason behind the general statement that entanglement is frame-independent Peres and Terno (2004). In Fucilla and Hatta (2026), it has been noticed that in the Regge limit,
@@
In the longitudinal case, the $q\bar{q}$ pair is always maximally entangled because $C_{nn}^{L}=1$ and $\Delta_{1}^{L}=2>0$. In the transverse case where $\Delta_{2}^{T}=-2C_{nn}^{T}\geq 0$, the pair is almost always entangled except when $C_{nn}^{T}=0$, which occurs in certain kinematical limits such as $k_{\perp}=0$ or $k_{\perp}=\infty$. The equation (57) is a direct consequence of the following relations obeyed by the components of $C^{L/T}$
@@
Away from the Regge limit, (58) continue to hold in the longitudinal case. However, in the transverse case we find from (41)-(III.2)
@@
Eq. (61) is an example of the so-called ‘Bell diagonal’ states Lang and Caves (2010) which are mixed states comprised only of the Bell states
@@
with $\sum_{i}p_{i}=1$. It is known that a Bell diagonal state is separable if $p_{i}\leq\frac{1}{2}$ for all $i$. In the present problem, the pair is separable only if $|X|=|Y|$ where $C_{nn}^{T}=0$. If either $|X|$ or $|Y|$ (accidentally) vanishes, then one of the Bell states is realized. Note that in the Regge limit, $X\approx(1-2z)I_{T}^{g}$ and $Y\approx I_{T}^{g}$. Thus the state is separable when $z=0,1$, and becomes the Bell state $|\Phi^{-}\rangle$ at $z=\frac{1}{2}$. Note that under Lorentz transformations, Bell states with sharp momenta remain Bell states up to local unitary transformations.
IV.2 Bell nonlocality
In terms of the correlation matrix $C$, the Bell-CHSH inequality reads Bell (1964); Clauser et al. (1969)
@@ IV.2 Bell nonlocality
where $\vec{n}_{i}$ ($i=1,2,3,4$) are unit vectors $|\vec{n}_{i}|=1$. Depending on the matrix $C_{ab}$, quantum two-qubit states can violate this inequality for deliberate choices of $\{\vec{n}_{i}\}$. When this happens, we say the $q\bar{q}$ pair exhibits ‘Bell nonlocality,’ and this property is frame-independent (see Barr et al. (2024) and references therein). The necessary and sufficient condition for Bell nonlocality is that the largest two of the three eigenvalues $\mu_{3}\leq\mu_{2}\leq\mu_{1}$ of the matrix $C^{\cal T}C$ satisfy Horodecki et al. (1995)
@@
Thus the inequality is always maximally violated in the longitudinal case $C_{nn}^{L}=1$, while in the transverse case it is violated if $C_{nn}^{T}\neq 0$. Since $C_{nn}^{T}\leq 0$, this happens to be the same condition as (55), (57) for nonvanishing entanglement Fucilla and Hatta (2026). In other words, Bell nonlocality and entanglement are equivalent. This is remarkable because, in general, the states that exhibit Bell nonlocality form a subset of the states that are entangled Werner (1989). Away from the Regge limit, the Bell-CHSH inequality continues to be maximally violated in the longitudinal case. In the transverse case, it is always violated for the $q=u,d$ quarks because (58) also holds for massless quarks. For massive quarks, there will be kinematical regions where the inequality is not violated. On general grounds, we expect that this region is broader than the region where the pair is separable (not entangled).
@@
IV.3 Magic
Finally, let us consider another measure of entanglement called ‘magic’. (See White and White (2024) for an introduction friendly to high energy physicists.) In quantum computing, it is known that entanglement alone does not guarantee quantum advantage Gottesman (1998). Entangled pairs with nonvanishing magic are crucial for quantum computers to outperform classical ones. Recently, this notion has been imported to the high energy physics community White and White (2024); Liu et al. (2025a); Gargalionis et al. (2025). One may ask, for example, how much magic is carried by elementary particle pairs produced in collider experiments. Of course, at present such a question is mostly of conceptual interest. But it may offer new perspectives on the nature of scattering in quantum field theories. To quantify magic, we adopt the ‘stabilizer Rényi entropy’ Leone et al. (2022); White and White (2024) which takes the following form for the present problem and in the present frame
@@ IV.3 Magic
Note that this quantity depends on the choice of frame and spinor basis Liu et al. (2025a). The following results should be understood as such.
@@
using the same basis as in (53). (The sign combination is arbitrary, except that the 14 and 41 components are complex conjugate to each other.) The result (71) is significantly smaller than the theoretical upper bound for pure two-qubit states Liu et al. (2025b)
@@
However, we point out that this limit can be saturated only by states with nonzero polarization vectors, namely, $B,\bar{B}\neq 0$ in the generic parametrization (26). Since $B^{L}=\bar{B}^{L}=0$, the maximal value (74) cannot be expected. A better comparison may be the range of $M_{2}$ for the subset of states with vanishing polarization. In Appendix A, we revisit the derivation of Liu et al. (2025b) by imposing an extra constraint $B=\bar{B}=0$ and obtain a stronger bound
@@
which is still larger than (71) by a factor of about 2. In Appendix A, we further show that (71) is the maximum value of $M_{2}$ if one restricts to an even smaller subset of states. We note that the numbers $\ln\frac{9}{5}$ and $\ln\frac{4}{3}$ have been previously encountered in Liu et al. (2025a) as the maximal amount of magic that can be generated from certain initial stabilizer states in QED processes.
V Numerical results
In this section we delineate kinematical regions in the parameter space $(z,k_{\perp})$ where the $q\bar{q}$ pair exhibits entanglement and Bell nonlocality. We also compute the stabilizer Rényi entropy $M_{2}^{T}$ for the transversely polarized photon. We employ the Goloskokov-Kroll model Goloskokov and Kroll (2007, 2008) of the quark and gluon GPDs. This model has been fitted to reproduce the exclusive vector meson production data at HERA. At small skewness, the gluon GPD features the Regge behavior
@@ V Numerical results
with $Q_{0}^{2}=4\,{\rm GeV}^{2}$ and $\mu_{h}=Q$ in electroproduction Goloskokov and Kroll (2007, 2008)). Thus the model should produce results consistent with Fucilla and Hatta (2026) in the Regge limit. Adapting to our process, we use $\mu_{h}={\rm max}\{Q,m,k_{\perp}\}$. The quark masses are fixed as $m_{u}=m_{d}=0$, $m_{s}=0.093$ GeV, $m_{c}=1.27$ GeV and $m_{b}=4.18$ GeV. For light quark production $q=u,d,s$, we include both the quark and gluon GPD contributions. For heavy quark production $q=c,b$, we include only the gluon GPD contribution (i.e., $I^{q}_{L/T}=0$ in this case) since the $c,b$-quark GPDs are unavailable (and expected to be negligible). Using the GK model, we numerically evaluate the integrals (10) on grid points on a $50\times 50$ lattice in the $(z,k_{\perp})$ plane
@@
and smoothly interpolate the results. $k_{\perp}^{\rm max}$ is chosen appropriately for a given value of the $\gamma+p$ CM energy $W$. The integrals can be evaluated in a numerically stable way. For very small values of $\xi$, this is achieved by using the contour deformation technique as explained around (43) of Hatta et al. (2025).
In electroproduction at $Q\neq 0$, the measured cross section (6) is the linear combination of the transverse and longitudinal contributions. Unfortunately, it is difficult to experimentally separate the two contributions at the EIC since typically $\varepsilon\approx 1$ at high energy colliders (see however Klest (2025)). This means that the density matrices must be averaged